OpenStudy (anonymous):

Need help with this problem:

OpenStudy (anonymous):

Write out the first four terms of the Maclaurin series of f(x): if f(0) = -2, f'(0) = -6. f''(0) = 7, f'''(0) = 13

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

@amistre64

ganeshie8 (ganeshie8):

Maclaurin series : \(\large f(x) = f(0) + xf'(0) + \frac{x^2}{2!}f''(0) + \frac{x^3}{3!}f'''(0) + ....\)

ganeshie8 (ganeshie8):

simply plugin the given values

OpenStudy (anonymous):

okay thanks!

ganeshie8 (ganeshie8):

u wlc !